# Non-universal surface magnetoelectric response in antiferromagnetic topological insulators

### v2

@@ Abstract
We employ Eq. (70) and (71,72) of Ref. [67] for $\overline{\bm{M}}^{(E)}$ and $\widetilde{\bm{M}}^{(E)}$, respectively, which applies to any crystalline insulator whose electronic ground state is characterized by a vanishing (triple of) Chern number(s). We will initially work in a general smooth (and periodic) Hamiltonian gauge, such that $U_{n\alpha}(\bm{k})\equiv\delta_{n\alpha}$ in those expressions; the WFs with respect to which the electronic contribution to the electric polarization and orbital magnetization are identified are |$\ket{W_{n,\bm{R}}}\propto\int_{\text{BZ}}d^{3}ke^{-i\bm{k}\cdot\bm{R}}\ket{\phi_{n,\bm{k}}}$. That is, we initially consider our gauge choice to be defined by unspecified Bloch energy eigenvectors |$\ket{\phi_{n,\bm{k}}}$ that are smooth over BZ and only later will we take |$\ket{\phi_{n,\bm{k}}}=\ket{\psi_{n,\bm{k}}}$ defined by Eq. (S40,S41). We denote the corresponding cell-periodic functions by |$\ket{v_{n,\bm{k}}}$ and identify a smooth global Hamiltonian gauge $\mathfrak{v}$ pointwise by $\mathfrak{v}_{\bm{k}}=(\ket{v_{n,\bm{k}}})_{n\in\{1,\ldots,N\}}$ where $N$ is the number of energy bands in the model. We choose to work in a Hamiltonian gauge because we seek the existence of a smooth global gauge in which Eq. (70) of Ref. [67] vanishes, and in any Hamiltonian gauge the terms involving $\mathcal{W}^{a}$ vanish. Otherwise, those terms appear to be generically nonzero. Perhaps in special cases smooth global symmetric gauge choices exist (although in a $\mathbb{Z}_{2}$-odd phase a time-reversal symmetric smooth global frame is topologically forbidden [68, 44]) for which these additional terms vanish. Nevertheless, we focus on smooth global Hamiltonian gauge choices. In such a gauge, Eq. (70) of Ref. [67] reduces to
@@
$+\frac{e^{2}}{2\hbar c}\epsilon^{iab}E^{l}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-2i\sum_{vv^{\prime}c}(\xi^{\mathfrak{v}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v}(\xi^{\mathfrak{v}})^{l}_{vc}+\sum_{vc}(\partial_{a}(\xi^{\mathfrak{v}})^{b}_{cv})(\xi^{\mathfrak{v}})^{l}_{vc}\Bigg{]},$ (S48)
where $\big{(}\xi^{\mathfrak{v}}\big{)}^{a}_{nm}(\bm{k})=i\innerproduct{v_{n,\bm{k}}}{\partial_{a}v_{m,\bm{k}}}$ are the components of the non-Abelian Berry connection induced by the gauge $\mathfrak{v}$. In what follows we will often not explicitly indicate the $\bm{k}$-dependence of quantities. Now recall that the explicit form of $\alpha^{il}_{\text{G}}$ is given by
@@
$\qquad\qquad\qquad\qquad\qquad\qquad-4\sum_{vcc^{\prime}}\frac{E_{c^{\prime},\bm{k}}-E_{c,\bm{k}}}{E_{v,\bm{k}}-E_{c,\bm{k}}}\text{Re}\big{[}i\xi^{a}_{cc^{\prime}}\xi^{b}_{c^{\prime}v}\xi^{l}_{vc}\big{]}\Bigg{)},$ (S49)
which is gauge invariant, hence the lack of a gauge label. In general, (generalized) TRS implies $\alpha^{il}_{\text{G}}=0$. In addition, note that in the second and third terms in the brackets of Eq. (S49), Berry connection matrix elements appear ($\xi^{a}_{vv^{\prime}}$ and $\xi^{a}_{cc^{\prime}}$, respectively) that can be independently varied by changing the crystal parameters of the Bloch Hamiltonian (while maintaining TRS). Thus, each of the terms in the brackets of Eq. (S49) must separately vanish if there is TRS. This implies that if the insulator exhibits TRS then all but the last line in the final equality in Eq. (S48) vanishes,
@@
$\equiv\bar{\alpha}_{\mathfrak{v}}^{li}E^{l}.$ (S50)
We can now implement any smooth global Hamiltonian gauge $\mathfrak{v}$ to evaluate Eq. (S50).
@@
$(\xi^{\mathfrak{v}})^{a}_{nm}(\bm{k})\equiv i\innerproduct{v_{n,\bm{k}}}{\partial_{a}v_{m,\bm{k}}}=\sum_{r,s=1}^{N}T^{\dagger}_{nr}(\bm{k})\big{(}(\xi^{\mathfrak{u}})^{a}_{rs}(\bm{k})+\mathcal{T}^{a}_{rs}(\bm{k})\big{)}T_{sm}(\bm{k}),$ (S52)
where $\mathcal{T}^{a}_{rs}(\bm{k})\equiv i\sum_{l=1}^{N}(\partial_{a}T_{rl}(\bm{k}))T^{\dagger}_{ls}(\bm{k})$ and $T^{\dagger}_{ls}(\bm{k})\equiv\big{(}T^{\dagger}(\bm{k})\big{)}_{ls}$. Note that since $T(\bm{k})$ is block-diagonal, so too is $\mathcal{T}^{a}(\bm{k})$. Thus, $(\xi^{\mathfrak{v}})^{a}_{vc}=\sum_{v^{\prime}c^{\prime}}T^{\dagger}_{v,v^{\prime}}(\xi^{\mathfrak{u}})^{a}_{v^{\prime}c^{\prime}}T_{c^{\prime}c}$. Plugging this into Eq. (S50) we find
@@
$=\bar{\alpha}_{\mathfrak{u}}^{li}+\frac{e^{2}}{2\hbar c}\epsilon^{iab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-i\sum_{vv^{\prime}c}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\mathcal{T}^{b}_{v^{\prime}v}(\xi^{\mathfrak{u}})^{l}_{vc}+i\sum_{vcc^{\prime}}\mathcal{T}^{a}_{cc^{\prime}}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (S53)
@@
S2.3.2 Specific considerations for our model
In past work [15] we studied the non-magnetic analog of the model employed here and considered whether there exists a gauge $\mathfrak{v}$ in which $\bar{\alpha}_{\mathfrak{v}}^{li}$ vanishes (independent of the choice of model parameters). That model has $N=4$ and we considered the initial electronic state of the material to be such that $M=2$ (i.e., $v\in\{1,2\}$ and $c\in\{3,4\}$). Consider general smooth global (and periodic) gauges $\mathfrak{u}$ and $\mathfrak{v}$ in the total Bloch bundle, the components of which are related pointwise as in Eq. (S51). We only consider gauges in the total Bloch bundle that decompose as a product of smooth gauges on the occupied and unoccupied subbundles thereof. For gauges of this type, in the non-magnetic model gauge transformations are represented by a $4\times 4$ block-diagonal unitary matrix consisting of two $2\times 2$ unitary blocks. The question is whether or not one can choose the two $2\times 2$ blocks of $T$ in such a way that $\mathcal{T}_{vv^{\prime}}^{a}$ and $\mathcal{T}_{cc^{\prime}}^{a}$ are such that the integrand of Eq. (S53) vanishes. (The integral as a whole may accidentally vanish in some gauge for a given set of material parameters due to fine tuning, but if the Hamiltonian parameters are slightly altered this fine tuning cancellation would be destroyed – it would be more general if the integrand itself always vanished.) Indeed in past work [15] we explicitly found a gauge in which this occurred.
Let’s now focus on the anti-ferromagnetic model. We first employ our particular gauge choice $\mathfrak{u}$ defined by the Bloch energy eigenvectors in Eq. (S40,S41). Considering the $i=l$ component of Eq. (S50) and using the relations (S44,S45), we find
@@
$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{v=3}^{4}\sum_{c=5}^{6}\Bigg{(}-i\sum_{v^{\prime}=3}^{4}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+i\sum_{c^{\prime}=5}^{6}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (S54)
In order for Eq. (S54) to generally vanish over the parameter space of the model, either the terms in each of the ()($\ldots$)( … ) must cancel one another or the two ()($\ldots$)( … ) terms must cancel each other. Since there is generally no relation between, for example, $(\xi^{\mathfrak{u}})^{l}_{17}$ and $(\xi^{\mathfrak{u}})^{l}_{35}$, the ()($\ldots$)( … ) terms cannot generally cancel each other, thus the second possibility is eliminated. The first possibility does not occur either, since there is generally no relationship between $(\xi^{\mathfrak{u}})^{a}_{12}$ and $(\xi^{\mathfrak{u}})^{a}_{78}$ in the first ()($\ldots$)( … ) of Eq. (S54) nor between $(\xi^{\mathfrak{u}})^{a}_{34}$ and $(\xi^{\mathfrak{u}})^{a}_{56}$ in the second ()($\ldots$)( … ). (In the original gauge $\mathfrak{u}$ employed in the non-magnetic case there was such a relationship between $(\xi^{\mathfrak{u}})^{a}_{v^{\prime}v}$ and $(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}$ for $v\neq v^{\prime}$ and $c\neq c^{\prime}$.)
@@
$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{v=3}^{4}\sum_{c=5}^{6}\Bigg{(}-i\sum_{v^{\prime}=3}^{4}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\big{(}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+\mathcal{T}^{b}_{v^{\prime}v}\big{)}+i\sum_{c^{\prime}=5}^{6}\big{(}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}+\mathcal{T}^{a}_{cc^{\prime}}\big{)}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (S55)
Again, the lack of a relation between, for example, $(\xi^{\mathfrak{u}})^{l}_{17}$ and $(\xi^{\mathfrak{u}})^{l}_{35}$, means that the two ()($\ldots$)( … ) terms cannot generally cancel one another. So for the integrand to vanish there must be cancellation within each ()($\ldots$)( … ). Using the relations $(\xi^{\mathfrak{u}})^{a}_{17}=-(\xi^{\mathfrak{u}})^{a}_{82}$ and $(\xi^{\mathfrak{u}})^{a}_{18}=(\xi^{\mathfrak{u}})^{a}_{72}$ given in (S44), the first ()($\ldots$)( … ) term in Eq. (S55) can be rewritten as
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{87}(\xi^{\mathfrak{u}})^{b}_{71}-2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{21}(\xi^{\mathfrak{u}})^{b}_{17}\bigg{)}\Bigg{]},$ (S56)
@@
Using the relations $(\xi^{\mathfrak{u}})^{a}_{35}=-(\xi^{\mathfrak{u}})^{a}_{64}$ and $(\xi^{\mathfrak{u}})^{a}_{36}=(\xi^{\mathfrak{u}})^{a}_{54}$, and $(\xi^{\mathfrak{u}})^{a}_{33}=-(\xi^{\mathfrak{u}})^{a}_{44}$ and $(\xi^{\mathfrak{u}})^{a}_{55}=-(\xi^{\mathfrak{u}})^{a}_{66}$ given in (S44), analogous arguments apply to the the second bracketed term in Eq. (S55). The sum total of these arguments is that there exists a smooth global Hamiltonian gauge $\mathfrak{v}$ in which $\bar{\alpha}^{ll}_{\mathfrak{v}}$ vanishes independent of material parameters only if there exists a $T$ relating smooth Hamiltonian gauges $\mathfrak{u}$ and $\mathfrak{v}$ that satisfies $(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{nm}=\delta_{nm}\partial_{a}\phi(\bm{k})$ (thus from Eq. (S52) we have $(\xi^{\mathfrak{v}})^{a}_{nm}=\delta_{nm}\partial_{a}\phi(\bm{k})$) for $n$ and $m$ in the same degenerate $2\times 2$ block of bands. And since $\mathfrak{v}$ is a Hamiltonian gauge (i.e., $T(\bm{k})$ and $\mathcal{T}^{a}(\bm{k})$ are block diagonal consisting of four $2\times 2$ blocks), Eq. (S52) implies that when $n$ and $m$ are not in the same degenerate $2\times 2$ block of bands, $(\xi^{\mathfrak{v}})^{a}_{nm}$ satisfies an analogous version of Eq. (S45); that is, $(\xi^{\mathfrak{v}})^{a}_{13},\ldots,(\xi^{\mathfrak{v}})^{a}_{86}$ vanishes unless $a=z$. If we assume that such a $T$ (and therefore $\mathfrak{v}$ related to $\mathfrak{u}$ by that $T$) exists, then
$\alpha^{\mathfrak{v}}_{\text{CS}}$ $=-\frac{e^{2}}{2\hbar c}\epsilon^{abd}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\left(\sum_{vv^{\prime}}(\xi^{\mathfrak{v}})^{a}_{vv^{\prime}}\partial_{b}(\xi^{\mathfrak{v}})^{d}_{v^{\prime}v}-\frac{2i}{3}\sum_{vv^{\prime}v_{1}}(\xi^{\mathfrak{v}})^{a}_{vv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v_{1}}(\xi^{\mathfrak{v}})^{d}_{v_{1}v}\right)=0,$ (S57)
where we have used $\epsilon^{abd}\partial_{b}\partial_{d}\phi=0$ and $\epsilon^{abd}(\partial_{b}\phi)(\partial_{d}\phi)=0$. This result is independent of material parameters. Indeed this is a contradiction since we have explicitly shown above that this model supports regions of parameter space in which the electronic ground state (at half filling) is $\mathbb{Z}_{2}$-odd and $\alpha^{\mathfrak{v}}_{\text{CS}}\text{ mod }\frac{e^{2}}{hc}=\frac{e^{2}}{2hc}$. Thus, there does not exist a smooth global Hamiltonian gauge $\mathfrak{v}$ in which $\bar{\alpha}^{ll}_{\mathfrak{v}}$ vanishes independent of material parameters. That is, in the bulk 3D AFM tight-binding model considered here, the magnetoelectric response cannot always be made entirely itinerant (with respect to a smooth global and periodic Hamiltonian gauge), in contrast to the non-magnetic case [15].

### v4

@@ B.3.1 Considerations for a general smooth global Hamiltonian gauge
We employ Eq. (70) and Eqs. (71) and (72) of Ref. [63] for $\overline{\bm{M}}^{(E)}$ and $\widetilde{\bm{M}}^{(E)}$, respectively, which applies to any crystalline insulator whose electronic ground state is characterized by a vanishing (triple of) Chern number(s). We will initially work in a general smooth (and periodic) Hamiltonian gauge, such that $U_{n\alpha}(\bm{k})\equiv\delta_{n\alpha}$ in those expressions; the WFs with respect to which the electronic contribution to the electric polarization and orbital magnetization are identified are |$\ket{W_{n,\bm{R}}}\propto\int_{\text{BZ}}d^{3}ke^{-i\bm{k}\cdot\bm{R}}\ket{\phi_{n,\bm{k}}}$. That is, we initially consider our gauge choice to be defined by unspecified Bloch energy eigenvectors |$\ket{\phi_{n,\bm{k}}}$ that are smooth over BZ and only later will we take |$\ket{\phi_{n,\bm{k}}}=\ket{\psi_{n,\bm{k}}}$ defined by Eqs. (49) and (50). We denote the corresponding cell-periodic functions by |$\ket{v_{n,\bm{k}}}$ and identify a smooth global Hamiltonian gauge $\mathfrak{v}$ pointwise by $\mathfrak{v}_{\bm{k}}=(\ket{v_{n,\bm{k}}})_{n\in\{1,\ldots,N\}}$ where $N$ is the number of energy bands in the model. We choose to work in a Hamiltonian gauge because we seek the existence of a smooth global gauge in which Eq. (70) of Ref. [63] vanishes, and in any Hamiltonian gauge the terms involving $\mathcal{W}^{a}$ vanish. Otherwise, those terms appear to be generically nonzero. Perhaps in special cases smooth global symmetric gauge choices exist (although in a $\mathbb{Z}_{2}$-odd phase a time-reversal symmetric smooth global frame is topologically forbidden [64, 44]) for which these additional terms vanish. Nevertheless, we focus on smooth global Hamiltonian gauge choices. In such a gauge, Eq. (70) of Ref. [63] reduces to
@@
$+\frac{e^{2}}{2\hbar c}\epsilon^{iab}E^{l}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-2i\sum_{vv^{\prime}c}(\xi^{\mathfrak{v}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v}(\xi^{\mathfrak{v}})^{l}_{vc}+\sum_{vc}(\partial_{a}(\xi^{\mathfrak{v}})^{b}_{cv})(\xi^{\mathfrak{v}})^{l}_{vc}\Bigg{]},$ (57)
where $\big{(}\xi^{\mathfrak{v}}\big{)}^{a}_{nm}(\bm{k})=i\innerproduct{v_{n,\bm{k}}}{\partial_{a}v_{m,\bm{k}}}$ are the components of the nonAbelian Berry connection induced by the gauge $\mathfrak{v}$. In what follows we will often not explicitly indicate the $\bm{k}$-dependence of quantities. Now recall that the explicit form of $\alpha^{il}_{\text{G}}$ is given by
@@
$\qquad\qquad\qquad\qquad\qquad\qquad-4\sum_{vcc^{\prime}}\frac{E_{c^{\prime},\bm{k}}-E_{c,\bm{k}}}{E_{v,\bm{k}}-E_{c,\bm{k}}}\text{Re}\big{[}i\xi^{a}_{cc^{\prime}}\xi^{b}_{c^{\prime}v}\xi^{l}_{vc}\big{]}\Bigg{)},$ (58)
which is gauge invariant, hence the lack of a gauge label. In general, (generalized) TRS implies $\alpha^{il}_{\text{G}}=0$. In addition, note that in the second and third terms in the brackets of Eq. (58), Berry connection matrix elements appear ($\xi^{a}_{vv^{\prime}}$ and $\xi^{a}_{cc^{\prime}}$, respectively) that can be independently varied by changing the crystal parameters of the Bloch Hamiltonian (while maintaining TRS). Thus, each of the terms in the brackets of Eq. (58) must separately vanish if there is TRS. This implies that if the insulator exhibits TRS then all but the last line in the final equality in Eq. (57) vanishes:
@@
$\equiv\bar{\alpha}_{\mathfrak{v}}^{li}E^{l}.$ (59)
We can now implement any smooth global Hamiltonian gauge $\mathfrak{v}$ to evaluate Eq. (59).
@@
$(\xi^{\mathfrak{v}})^{a}_{nm}(\bm{k})\equiv i\innerproduct{v_{n,\bm{k}}}{\partial_{a}v_{m,\bm{k}}}=\sum_{r,s=1}^{N}T^{\dagger}_{nr}(\bm{k})\big{(}(\xi^{\mathfrak{u}})^{a}_{rs}(\bm{k})+\mathcal{T}^{a}_{rs}(\bm{k})\big{)}T_{sm}(\bm{k}),$ (61)
where $\mathcal{T}^{a}_{rs}(\bm{k})\equiv i\sum_{l=1}^{N}(\partial_{a}T_{rl}(\bm{k}))T^{\dagger}_{ls}(\bm{k})$ and $T^{\dagger}_{ls}(\bm{k})\equiv\big{(}T^{\dagger}(\bm{k})\big{)}_{ls}$. Note that since $T(\bm{k})$ is block-diagonal, so too is $\mathcal{T}^{a}(\bm{k})$. Thus, $(\xi^{\mathfrak{v}})^{a}_{vc}=\sum_{v^{\prime}c^{\prime}}T^{\dagger}_{v,v^{\prime}}(\xi^{\mathfrak{u}})^{a}_{v^{\prime}c^{\prime}}T_{c^{\prime}c}$. Plugging this into Eq. (59) we find
@@
$=\bar{\alpha}_{\mathfrak{u}}^{li}+\frac{e^{2}}{2\hbar c}\epsilon^{iab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-i\sum_{vv^{\prime}c}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\mathcal{T}^{b}_{v^{\prime}v}(\xi^{\mathfrak{u}})^{l}_{vc}+i\sum_{vcc^{\prime}}\mathcal{T}^{a}_{cc^{\prime}}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (62)
@@
B.3.2 Specific considerations for our model
In past work [15] we studied the nonmagnetic analog of the model employed here and considered whether there exists a gauge $\mathfrak{v}$ in which $\bar{\alpha}_{\mathfrak{v}}^{li}$ vanishes (independent of the choice of model parameters). That model has $N=4$ and we considered the initial electronic state of the material to be such that $M=2$ (i.e., $v\in\{1,2\}$ and $c\in\{3,4\}$). Consider general smooth global (and periodic) gauges $\mathfrak{u}$ and $\mathfrak{v}$ in the total Bloch bundle, the components of which are related pointwise as in Eq. (60). We only consider gauges in the total Bloch bundle that decompose as a product of smooth gauges on the occupied and unoccupied subbundles thereof. For gauges of this type, in the nonmagnetic model gauge transformations are represented by a $4\times 4$ block-diagonal unitary matrix consisting of two $2\times 2$ unitary blocks. The question is whether or not one can choose the two $2\times 2$ blocks of $T$ in such a way that $\mathcal{T}_{vv^{\prime}}^{a}$ and $\mathcal{T}_{cc^{\prime}}^{a}$ are such that the integrand of Eq. (62) vanishes. (The integral as a whole may accidentally vanish in some gauge for a given set of material parameters due to fine tuning, but if the Hamiltonian parameters are slightly altered this fine tuning cancellation would be destroyed – it would be more general if the integrand itself always vanished.) Indeed in past work [15] we explicitly found a gauge in which this occurred.
Let us now focus on the antiferromagnetic model. We first employ our particular gauge choice $\mathfrak{u}$ defined by the Bloch energy eigenvectors in Eqs. (49) and (50). Considering the $i=l$ component of Eq. (59) and using the relations (53) and (54), we find
@@ B.3.2 Specific considerations for our model
$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{v=3}^{4}\sum_{c=5}^{6}\Bigg{(}-i\sum_{v^{\prime}=3}^{4}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+i\sum_{c^{\prime}=5}^{6}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (63)
For Eq. (63) to generally vanish over the parameter space of the model, either the terms in each of the ()($\ldots$)( … ) must cancel one another or the two ()($\ldots$)( … ) terms must cancel each other. Since there is generally no relation between, for example, $(\xi^{\mathfrak{u}})^{l}_{17}$ and $(\xi^{\mathfrak{u}})^{l}_{35}$, the ()($\ldots$)( … ) terms cannot generally cancel each other, thus the second possibility is eliminated. The first possibility does not occur either, since there is generally no relationship between $(\xi^{\mathfrak{u}})^{a}_{12}$ and $(\xi^{\mathfrak{u}})^{a}_{78}$ in the first ()($\ldots$)( … ) of Eq. (63) nor between $(\xi^{\mathfrak{u}})^{a}_{34}$ and $(\xi^{\mathfrak{u}})^{a}_{56}$ in the second ()($\ldots$)( … ). (In the original gauge $\mathfrak{u}$ employed in the nonmagnetic case there was such a relationship between $(\xi^{\mathfrak{u}})^{a}_{v^{\prime}v}$ and $(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}$ for $v\neq v^{\prime}$ and $c\neq c^{\prime}$.)
@@
$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{v=3}^{4}\sum_{c=5}^{6}\Bigg{(}-i\sum_{v^{\prime}=3}^{4}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\big{(}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+\mathcal{T}^{b}_{v^{\prime}v}\big{)}+i\sum_{c^{\prime}=5}^{6}\big{(}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}+\mathcal{T}^{a}_{cc^{\prime}}\big{)}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (64)
Again, the lack of a relation between, for example, $(\xi^{\mathfrak{u}})^{l}_{17}$ and $(\xi^{\mathfrak{u}})^{l}_{35}$, means that the two ()($\ldots$)( … ) terms cannot generally cancel one another. So for the integrand to vanish there must be cancellation within each ()($\ldots$)( … ). Using the relations $(\xi^{\mathfrak{u}})^{a}_{17}=-(\xi^{\mathfrak{u}})^{a}_{82}$ and $(\xi^{\mathfrak{u}})^{a}_{18}=(\xi^{\mathfrak{u}})^{a}_{72}$ given in (53), the first ()($\ldots$)( … ) term in Eq. (64) can be rewritten as
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{87}(\xi^{\mathfrak{u}})^{b}_{71}-2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{21}(\xi^{\mathfrak{u}})^{b}_{17}\bigg{)}\Bigg{]},$ (65)
@@
Using the relations $(\xi^{\mathfrak{u}})^{a}_{35}=-(\xi^{\mathfrak{u}})^{a}_{64}$ and $(\xi^{\mathfrak{u}})^{a}_{36}=(\xi^{\mathfrak{u}})^{a}_{54}$, and $(\xi^{\mathfrak{u}})^{a}_{33}=-(\xi^{\mathfrak{u}})^{a}_{44}$ and $(\xi^{\mathfrak{u}})^{a}_{55}=-(\xi^{\mathfrak{u}})^{a}_{66}$ given in Eq. (53), analogous arguments apply to the the second bracketed term in Eq. (64). The sum total of these arguments is that there exists a smooth global Hamiltonian gauge $\mathfrak{v}$ in which $\bar{\alpha}^{ll}_{\mathfrak{v}}$ vanishes independent of material parameters only if there exists a $T$ relating smooth Hamiltonian gauges $\mathfrak{u}$ and $\mathfrak{v}$ that satisfies $(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{nm}=\delta_{nm}\partial_{a}\phi(\bm{k})$ (thus from Eq. (61) we have $(\xi^{\mathfrak{v}})^{a}_{nm}=\delta_{nm}\partial_{a}\phi(\bm{k})$) for $n$ and $m$ in the same degenerate $2\times 2$ block of bands. And since $\mathfrak{v}$ is a Hamiltonian gauge (i.e., $T(\bm{k})$ and $\mathcal{T}^{a}(\bm{k})$ are block diagonal consisting of four $2\times 2$ blocks), Eq. (61) implies that when $n$ and $m$ are not in the same degenerate $2\times 2$ block of bands, $(\xi^{\mathfrak{v}})^{a}_{nm}$ satisfies an analogous version of Eq. (54); that is, $(\xi^{\mathfrak{v}})^{a}_{13},\ldots,(\xi^{\mathfrak{v}})^{a}_{86}$ vanishes unless $a=z$. If we assume that such a $T$ (and therefore $\mathfrak{v}$ related to $\mathfrak{u}$ by that $T$) exists, then
$\alpha^{\mathfrak{v}}_{\text{CS}}$ $=-\frac{e^{2}}{2\hbar c}\epsilon^{abd}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\left(\sum_{vv^{\prime}}(\xi^{\mathfrak{v}})^{a}_{vv^{\prime}}\partial_{b}(\xi^{\mathfrak{v}})^{d}_{v^{\prime}v}-\frac{2i}{3}\sum_{vv^{\prime}v_{1}}(\xi^{\mathfrak{v}})^{a}_{vv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v_{1}}(\xi^{\mathfrak{v}})^{d}_{v_{1}v}\right)=0,$ (66)
where we have used $\epsilon^{abd}\partial_{b}\partial_{d}\phi=0$ and $\epsilon^{abd}(\partial_{b}\phi)(\partial_{d}\phi)=0$. This result is independent of material parameters. Indeed this is a contradiction since we have explicitly shown above that this model supports regions of parameter space in which the electronic ground state (at half filling) is $\mathbb{Z}_{2}$-odd and $\alpha^{\mathfrak{v}}_{\text{CS}}\text{ mod }\frac{e^{2}}{hc}=\frac{e^{2}}{2hc}$. Thus, there does not exist a smooth global Hamiltonian gauge $\mathfrak{v}$ in which $\bar{\alpha}^{ll}_{\mathfrak{v}}$ vanishes independent of material parameters. That is, in the bulk 3D AFM tight-binding model considered here, the magnetoelectric response cannot always be made entirely itinerant (with respect to a smooth global and periodic Hamiltonian gauge), in contrast to the nonmagnetic case [15].